Question 3 (1 point)

Fill in the blank. Speed Car Rental company found that the tire lifetime for their vehicles has a mean of 49,038.99 miles and standard deviation of 5,415.52 miles. In a sample of 69 tires tested, there is a 78% chance that the average tire lifetime is less than ________ miles.

Question 3 options:

1)

44,857.16

2)

There is not enough information to determine this.

3)

53,220.82

4)

49,542.42

5)

48,535.56
Answer #1

solution

mean = = 49038.99

standard deviation = = 5415.52

Using standard normal table,

P(Z < z) =78 %

= P(Z < z) = 0.78  

= P(Z <0.77 ) = 0.78

z = 0.77

Using z-score formula  

x= z * +

x= 0.77*5415.52+49038.99

x= 53220.82